OpenStudy (anonymous):

Write the equation of a line in POINT SLOPE FORM that has a slope of -3/4 and passes through the point (4,5)

OpenStudy (anonymous):

plug in to the formula: y - y1 = m(x - x1) where m is your slope, and (x1, y1) is the point the line passes through.

OpenStudy (anonymous):

dpainc gave the approach.

OpenStudy (karatechopper):

point slope form...do you know what that is kaytree?

OpenStudy (karatechopper):

like the formula?

OpenStudy (anonymous):

@karatechopper i learned some of it but not all

OpenStudy (karatechopper):

@Kaytree16 ok well lets walk you through this if you don't know, is that fine?

OpenStudy (anonymous):

@karatechopper , thanks(:

OpenStudy (karatechopper):

answers won't get you anywhere, UNDERSTANDING gets you everywhere

OpenStudy (karatechopper):

@luisrock2008 ^

OpenStudy (karatechopper):

alrighty kaytree! lets begin with figuring out what point slope form is. the point slope form formula is y-y1=m(x-x1)

OpenStudy (karatechopper):

do you know what the y1 and x1 are?

OpenStudy (karatechopper):

now now please do not be greedy and mean:) we want a clean enviorment:D

OpenStudy (karatechopper):

@Kaytree16 ^^ look at what i said please:)

OpenStudy (anonymous):

@karatechopper , yes

OpenStudy (karatechopper):

alright good! can i know what grade you are in so that i can make my explanation better?

OpenStudy (anonymous):

Sophmore

OpenStudy (karatechopper):

ok well im a 7th grader so please excuse me if i sound childish:)

OpenStudy (anonymous):

lol, your a smart 7th Grader .

OpenStudy (karatechopper):

thank u!

OpenStudy (karatechopper):

ok well getting back to this problem..

OpenStudy (karatechopper):

first things first, do you know what your slope is?

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

it says so right up there in your question tho...

OpenStudy (anonymous):

Oh, sorry -3/4

OpenStudy (karatechopper):

:) much better;) ok so do you know which variable the slope takes the place of in that formula?

OpenStudy (anonymous):

... (4,5) ?

OpenStudy (karatechopper):

lol nope! the slope is always the variable "m". therefore you would put the -3/4 in the place of the m in the point slope form equation

OpenStudy (anonymous):

ohh , Okay lol

OpenStudy (anonymous):

why was my answer deleted?

OpenStudy (karatechopper):

haha yea:P ok so to get a better inderstanding lets rewrite the point slope form formula and instead of placing an m there, we will place the actual slope for this problem!

OpenStudy (karatechopper):

@FoolForMath deleted it:)

OpenStudy (anonymous):

why? I gave her the right answer

OpenStudy (karatechopper):

we will see if it is correct when me and @Kaytree16 r done solving it

OpenStudy (anonymous):

^ @karatechopper thank you(:

OpenStudy (karatechopper):

No problem!

OpenStudy (karatechopper):

back to the problem tho..can you rewrite the point slope form for me @Kaytree16

OpenStudy (anonymous):

@luisrock2008: Your answer has almost no explanation with it so it was not helping the OP. Besides that your answer was *incorrect*.

OpenStudy (karatechopper):

whats OP?

OpenStudy (anonymous):

OP=Original poster

OpenStudy (karatechopper):

ahh

OpenStudy (anonymous):

@FoolForMath can you just make him stop commenting .hes not helping .

OpenStudy (anonymous):

m=-3/4 ?

OpenStudy (anonymous):

No but I used the equation @dpalnc gave us. I dein't even get a medal for trying. I also din't know we needed to show work

OpenStudy (anonymous):

He/She is already warned once. If he continue to do this, I will have to suspend him/her.

OpenStudy (karatechopper):

yes, now write it in the point slope form formula

OpenStudy (anonymous):

... How

OpenStudy (anonymous):

im confused again

OpenStudy (anonymous):

@FoolForMath thank you

OpenStudy (karatechopper):

by doing this! y-y1=m(x-x1) y-y1=-3/4(x-x1)

OpenStudy (anonymous):

ahhhh, okay

OpenStudy (karatechopper):

yep! now it comes to that one point of (4,5)

OpenStudy (anonymous):

Okay

OpenStudy (karatechopper):

ok so you told me you know what the x1 and y1 is. so this should be easy for you to plug the (4,5) in the equation if you r a sophmore;)

OpenStudy (anonymous):

i said i know what they are not how to .

OpenStudy (karatechopper):

yes yes, but that should help u! x1 is the x coordinate and y1 is the y coordinate therefore you just put the x value in for the x and the y value in for the y

OpenStudy (anonymous):

ahhhhhh i see

OpenStudy (karatechopper):

yep so would u please put those values in with out new formula? y-y1=-3/4(x-x1)

OpenStudy (anonymous):

5-5=-3/4(4-4) ?

OpenStudy (karatechopper):

u only plug it in for the x1 and y1 not the x and y

OpenStudy (anonymous):

ohh, /:

OpenStudy (karatechopper):

yea i know math can get confusing! dont stress!

OpenStudy (anonymous):

imm not stressed just confused again /:

OpenStudy (anonymous):

where did you go ?/:

OpenStudy (anonymous):

y-51=-3/4(x-41)????

OpenStudy (anonymous):

@FoolForMath could you help?

OpenStudy (anonymous):

easiest way is to plug in the numbers on this y-y1=m(x-x1)

OpenStudy (anonymous):

i did .

OpenStudy (anonymous):

here is step 1: y-5= -3/4 (x - 4)

OpenStudy (anonymous):

step 2 is to multiply the slope by the numbers in the bracket y-5 = -3/4x -3

OpenStudy (anonymous):

then you need to get rid of 5 so you add 5 on both sides therefore. -5+5 = 0 and -3+5=2

OpenStudy (anonymous):

answer is y = -3/4 +2

OpenStudy (anonymous):

thank you .

OpenStudy (anonymous):

can i have a medal.

OpenStudy (anonymous):

idk how